OpenStudy (anonymous):

Solve :

OpenStudy (anonymous):

\[\frac{ 2x }{ 3x-1 }= \frac{ 3 }{ 4 }\]

OpenStudy (anonymous):

first multiply both equations by 3x-1 and 4

OpenStudy (anonymous):

2x(4)=3(3x-1)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{ 2x }{ 3x-1} - \frac{ 3 }{4 }=0\]

OpenStudy (anonymous):

no i have to make it equal zero first

OpenStudy (anonymous):

@burhan101 that method is much more complicated than it needs to be

OpenStudy (anonymous):

i agree :) but it's the way my teacher would like me to do it

OpenStudy (anonymous):

well...ok then

OpenStudy (anonymous):